OpenStudy (anonymous):

The equations of two lines are x - 3y = 6 and y = 3x + 2. Determine if the lines are parallel, perpendicular or neither, i dont know how to find out if its parallel or perpendicular

OpenStudy (jdoe0001):

what's the slope of the 2nd one? that is y = 3x + 2.

OpenStudy (anonymous):

3x

OpenStudy (jdoe0001):

hmm have you covered slopes yet?

OpenStudy (anonymous):

yes, im just having a bit of a hard time with it

OpenStudy (jdoe0001):

hmmm anyhow.... the slope would be 3 only without the "x" so the idea being.. parallel - they both have the same exact slope perpendicular - the slope of one, is the same as the other BUT upside down and negative neither - see above =) so... the let us see their slopes then \(bf x-3y=6\implies -3y=-x+6\implies y=\cfrac{-x+6}{-3}\implies y=\cfrac{x}{-3}+\cfrac{\cancel{ 6 }}{\cancel{ -3 }} \\ \quad \\ y={\color{brown}{ -\cfrac{1}{3}}}x+2\) and the slope of the 2nd one is \(\bf y = {\color{brown}{ 3}}x + 2\) now, look at the slopes..... what do you think? paralle, perpendicular or neither?

OpenStudy (anonymous):

neither?

OpenStudy (jdoe0001):

hmm what makes you think they're not parallel? or perpendicular?

OpenStudy (anonymous):

the slopes arent the same and they arent negative

OpenStudy (anonymous):

but the y intercepts are the same

OpenStudy (jdoe0001):

hmm notice the "dash" in front of the 1st slope, that one is negative

OpenStudy (jdoe0001):

so they're not the same..... true but is one the same as the other upside down and negative?

OpenStudy (anonymous):

OOOOOOOHHHH!!!! I GET IT

OpenStudy (jdoe0001):

keep in mind that anynumber is the same as anynumber/1 that is \(\bf 5=\cfrac{5}{1}\qquad 1,000=\cfrac{1,000}{1}\) and so on

OpenStudy (anonymous):

okay, thank you so much for your help :3

OpenStudy (jdoe0001):

\(\bf x-3y=6\implies -3y=-x+6\implies y=\cfrac{-x+6}{-3}\implies y=\cfrac{x}{-3}+\cfrac{\cancel{ 6 }}{\cancel{ -3 }} \\ \quad \\ y={\color{brown}{ -\cfrac{1}{3}}}x+2 \\ \quad \\ \quad \\ y = {\color{brown}{ 3}}x + 2\implies y = {\color{brown}{ \cfrac{+3}{1}}}x + 2\)

OpenStudy (jdoe0001):

yw